order: Potassium Chloride 10mEq/hr IV for arrhythmias

Available: 2 mEq/mL

How many milliliters will the client receive each day?
Answer #1

Amount of potassium to be infused per hour= 10 mEq/hr

Available - 2 mEq/ml

So, volume of potassium chloride solution to be infused per hour = 10/2 = 5ml.

Total volume of potassium to be recieved in one day = 5*24 = 120 ml.